# Automated System-level Testing of Unmanned Aerial Systems

### v1

@@ 3.3.3 Specifying Expected Behavior of the SUT
After modeling the domain UAS SUT and the flight behavior of the UAV, the expected behavior of UAS SUT in the form of OCL constraints is required to be specified. The OCL constraints should contain the expected ranges of value on various properties of the UAS SUT and w.r.t. the flight states. The purpose of the constraints is to check the invalid property values during the UAV flight. When formulating OCL constraints corresponding to UAV flight states, the specified ranges for a particular state must be consistent and non-overlapping with the subsequent state. For example, if the maximum altitude limit for Takeoff state is 50 feet, the altitude range for the subsequent state (Ascend) should consider 50 feet as the lower bound. Furthermore, the ranges defined in OCL constraints associated with UAV flight states should align with those specified in the general constraints. The outcomes of the OCL constraints evaluation play a substantial role in calculating the reward and validating the expected behavior of the UAV. Therefore, the constraints need to be carefully modeled to capture the required (valid) range of values for the different properties during different flight states. Listing 1 shows ten OCL constraints as an example. Among those constraints (C1& C2), two are general which means these constraints must be true regardless of any flight state. The remaining eight constraints (C3C10) are state invariants modeled for the Takeoff, Landing, Ascend, and Descend states. Constraint C1 models the valid altitude values range with 300m maximum altitude for the ArduCopter. Constraint C2 shows the valid values range for the distance in which the maximum distance the ArduCopter can cover is 5000m. Constraints C3 and C4 capture the valid range for the thrust and altitude_AGL during the Takeoff state. Similarly, constraints C5 and C6 capture the valid range of values for the airspeed and speed along the z-axis during the Landing state. Constraints C7 and C8 define valid ranges for airspeed (i.e., between 0 and 100 knots) and groundspeed (i.e., between 0 and 10 knots) during the Ascend state. Moreover, constraints C9 and C10 specify valid ranges for airspeed (specifically between 0 and 100 knots) and altitude_AGL (specifically between 10 and 100 feet) during the Descend state.
@@ 3.4 System-level Testing of UAS
A typical practice of the avionics tester is to simulate the flight behavior of the UAV using the SIL/HIL simulator. During the simulation, the tester performs a certain action, observes the next state of the UAV in an environment, and analyzes the flight data to see if each avionics component is behaving according to the specification. This process of testing requires control of the UAV, continuously deciding various actions, and observing state change in the UAV and environment. To automate this process, human-level decision-making and control of the UAV are required at runtime. Such a problem can be devised in the form of a Markov Decision Process (MDP) [23]. In this case, an AI technique, deep reinforcement learning (DRL) is an appropriate choice as compared to meta-heuristics search algorithms. Moreover, DRL provides a mechanism for on-the-fly testing of UAS in a dynamic environment which cannot be achieved by traditional model-based testing approaches (e.g., [24]). Using DRL, our approach introduces an AITester in place of a human tester to perform the system-level testing of the UAV. We formulate this problem as an MDP which consists of a set of states, a set of possible actions that can be performed in a state, a transition model, and a reward function [23]. For this purpose, the proposed profiles, UAS structural profile, and UAV behavioral profile. play an important role. More discussion on the adaptation of DRL to automate the system-level testing of the UAV is given in the following subsections.
@@ 3.4.1 Definitions
Definition 1 (State Space). A state space is defined as a ordered set of states $S=\{S_{0},S_{1},S_{2},\dots,S_{g}\}$ in which each state is represented by a tuple $S_{x}={\left\langle s_{0},s_{1},\dots,s_{9}\right\rangle}$. Where $S_{g}$ represents the goal (final) state and $s_{0}$ represents one of the flight state stereotypes modeled in the UAV flight state machine. The remaining state elements are denoted as $s_{1}:=altitude$, $s_{2}:=airspeed$, $s_{3}:=groundspeed$, $s_{4}:=roll$, $s_{5}:=pitch$, $s_{6}:=yaw$, $s_{7}:=heading$, $s_{8}:=battery$, and $s_{9}:=distance$.
Fig. 9 shows an example of the different states in the state space considering the UAV flight state machine shown in Fig. 7. For the given example, the initial state is represented as
[…]
@@
Definition 5 (Observation Space). An observation space is defined as a vector of UAV properties, i.e., $O=[o_{0},o_{1},o_{2},\dots,o_{n}]$ from the UAS SUT domain model created using the UAS structural profile. Where $o_{x}$ corresponds to the UAV properties such as thrust, airspeed, vx, vy, and vz.
@@ 3.4.2 Reward Function
The objective AITester is to take control of the UAV flight during simulation, perform various actions on the UAV, and maximize the number of violations of the expected behavior. The reward function is designed based on two parameters, (i) correct/incorrect actions, and (ii) violations of the expected behavior specified as OCL constraints. These two parameters are important because a series of correct/incorrect actions may lead a UAV to crash. For example, if AITester continues to execute a correct action decreaseAlt() multiple times in Descent state, the UAV can hit the ground leading it to crash. This can happen frequently during training because the inherent behavior of the DRL algorithm involves exploring various possibilities of actions and observing their outcomes. Therefore, a correct action leading the UAV to crash is not considered a fault in the UAV. To handle such a situation, our approach utilizes the violation of the expected behavior of a UAV specified in the form of OCL constraints. The trace of violations in expected behavior can help in diagnosing faults in the SUT. Since OCL constraints are specified corresponding to each flight state of the UAV, it is possible to have multiple violations of OCL constraints at different states during the whole flight.
@@
$r=\left\{\begin{array}[]{r}1+m,\ \ correct\ \ action\\ -1,\ \ incorrect\ \ action\end{array}\right.$ (2)
The cumulative reward for a single episode is calculated using Equation 3. According to this Equation, the cumulative reward is the added sum of the reward for individual action performed at time t (calculated using Equation 2) multiplied with discount factor ${\gamma}^{t}$ at time step t. The values of the discount factor $\gamma$ are between 0 and 1, i.e., $\gamma\in[0,1]$. The discount factor determines the AITester’s preference between the current and future rewards. If the value of $\gamma$ is close to 0, the AITester will prefer current rewards over future rewards. If the value of $\gamma$ is close to 1, the AITester will prefer future rewards over the current rewards. The purpose of factor ${\gamma}^{t}$ is to set a value close to 1 and progressively decrease the value after each time step t. In this way, AITester gives preference to future rewards at initial time steps and moves slowly towards current rewards.
@@
3.4.3 Long Short-Term Memory (LSTM) Network
The UAV flight data at a particular time is correlated to the flight data at the previous instance of time. Due to this UAV flight data holds a temporal relationship based on environmental context which is highly dynamic. Retaining information among these data points in subsequent time steps is important to comprehend the flight context. Moreover, our approach utilizes Deep Q-Network (DQN) [25] which works on the experience replay method by taking random samples from the replay memory of past experiences. This is an inherent feature of DQN to make learning efficient, however, it breaks correlation among different data points which makes it difficult to acquire the flight context. In such a scenario, among different types of artificial neural networks, the recurrent neural network (RNN) specifically the LSTM network architecture is most suitable for the temporal nature of data and capturing the state context. The LSTM network consists of one or more input gates, output gates, forget gates, and memory cells. The core purpose of the memory cell is to capture the correlation between the current flight data at time t with preceding flight data points [i.e., (t-1)!].
Figure 10: A single layer representation of LSTM architecture used with the RL algorithm
@@ 4 Implementation
We implement the proposed approach in a toolset to support the automated system-level testing of the UAV. All main modules of the toolset are packed into two separate packages. The toolset is available at the GitHub repository1. In the following subsections, we discuss every module of the toolset w.r.t. each package.
@@ 4.1.1 Model Loader
First, this module takes the UAS SUT domain model in the form of UML and loads this model in a data structure along with the meta-class elements (i.e., profile stereotypes). Second, it takes the flight behavior model in the form of a UML state machine and traverses the model to extract flight states, events, and transitions. In case the state machine has composite and orthogonal states, these states are flattened using the algorithm presented in [24]. For the modeling manipulation, it uses an external API named Eclipse Modeling Framework (EMF [^3] https://www.eclipse.org/modeling/emf/) that allows it to load and read the model information conveniently. The third function of this module is to load the OCL constraints and state invariants.
@@ 4.1.4 Constraints Evaluator
The core functionality of this module is to evaluate the OCL constraints and state invariants on the instance model containing the UAV flight information. For this purpose, the Eclipse OCL [^4] https://projects.eclipse.org/projects/modeling.mdt.ocl evaluator is used. Before starting the evaluation process, the OCL constraints and state invariants are extracted (using Model Loader) corresponding to each state of the UAV flight. During execution, the flight data received from the environment is sent to Instance Populator. The instance model is used to prepare an OCL evaluator environment for the evaluation of OCL constraints [29, 30]. Before starting the evaluation process, OCL constraints are loaded from a single file corresponding to each state of the UAV flight. This file contains general constraints and state-specific constraints, resembling the format depicted in Listing 1. All OCL constraints and state invariants are evaluated on the populated instance model. If an OCL constraint or state invariant fails on the instance model, the Eclipse OCL evaluator returns false. If the OCL constraint or state invariant is passed on the instance model, the Eclipse OCL evaluator returns true. At the end of the constraints evaluation, the evaluator returns the total number of failed OCL constraints (violations in the expected behavior of the SUT) for the given flight data.
@@ 4.2 AITester
This is the main part of the toolset. The key functions of this component are UAV environment preparation, interaction with Constraints Evaluator for the reward calculation, and training and evaluation modes for system-level testing of UAVs. Since AITester is aimed to be used at different levels of testing, the tool contains a training module as well as an evaluation module. All modules of the AITester are developed using Python programming language. The deep learning framework used for this purpose is PyTorch [31].
@@ 5 Empirical Evaluation
This section presents the empirical evaluation of AITester including two experiments. The first experiment is performed using the ArduCopter autopilot system and the second experiment is performed using an industrial case study of ground control station (GCS) cockpit display systems (CDS), i.e., GCS-CDS. In the following subsections, we discuss research questions, the details of both case studies, the setup of both experiments, the design and execution of the experiments, the results of both experiments, and the possible threats to the validity of experimental results.
@@ 5.1 Research Questions
AITester utilizes a UAV flight state machine for the system-level testing of UAS system under test (SUT) with the goal of violating the expected behavior of the SUT. There are two key aspects for analyzing the performance of an AITester. The first is the effectiveness of AITester in generating test scenarios that can lead to deviations from the expected behavior of the UAS SUT. During the process of testing, the second aspect is the ability of the AITester to explore diverse paths from the UAV flight state machine. We formulate research questions for the experimental evaluation based on these two aspects.
@@
Both RQs are important from the testing perspective of a UAS SUT. The goal of the AITester is to perform system-level testing of a UAS SUT using a state machine representing the flight behavior of a UAV. RQ1 assesses AITester in terms of violations of the expected behavior of a UAS SUT that can lead to potential faults. This is a primary concern of testing a UAS. Since a flight state machine is utilized in the testing process, it is quite possible that AITester learns to explore a specific set of flight states and actions leading to maximum violations using the same set of OCL constraints. This results in testing of the same behavior multiple times which is not a desired feature. Therefore, the analysis of diversity in exploring flight paths is an important consideration.

### v2

@@ 3.3.3 Specifying Expected Behavior of the SUT
After modeling the domain UAS SUT and the flight behavior of the UAV, the expected behavior of UAS SUT in the form of OCL constraints is required to be specified. The OCL constraints should contain the expected ranges of value on various properties of the UAS SUT and w.r.t. the flight states. The purpose of the constraints is to check the invalid property values during the UAV flight. When formulating OCL constraints corresponding to UAV flight states, the specified ranges for a particular state must be consistent and non-overlapping with the subsequent state. For example, if the maximum altitude limit for Takeoff state is 50 feet, the altitude range for the subsequent state (Ascend) should consider 50 feet as the lower bound. Furthermore, the ranges defined in OCL constraints associated with UAV flight states should align with those specified in the general constraints.
Listing 1: OCL constraints including the general and the state invariants
--altitude values range (in meters)
C1: context UAV inv: self.location.altitude_AGL>0 and self.location.altitude_AGL<=300
--distance values range (in meters)
C2: context UAV inv: self.rangefinder.distance>0 and self.rangefinder.distance<=5000
--State invariants for Takeoff
C3: context UAV inv: self.oclIsInState(Takeoff) and self.thrust>0 and self.thrust<1
C4: context UAV inv: self.oclIsInState(Takeoff) and self.location.altitude_AGL>0 and self.location.altitude_AGL<=50
--State invariants for Landing
C5: context UAV inv: self.oclIsInState(Landing) and self.airspeed>0 and self.airspeed<=5
C6: context UAV inv: self.oclIsInState(Landing) and self.speed.vz>0 and self.speed.vz<=5
--State invariants for Ascend
C7: context UAV inv: self.oclIsInState(Ascend) and self.airspeed>10 and self.airspeed<100
C8: context UAV inv: self.oclIsInState(Ascend) and self.groundspeed>0 and self.groundspeed<10
--State invariants for Descent
C9: context UAV inv: self.oclIsInState(Descent) and self.airspeed>=5 and self.airspeed<100
C10: context UAV inv: self.oclIsInState(Descent) and self.location.altitude_AGL>10 and self.location.altitude_AGL<100
OCL constraints can be derived from two key sources: (i) the UAV configuration manual, and (ii) domain experts or testers who define constraints based on specific testing contexts. UAV configuration manuals usually specify parameter ranges for general scenarios and specific flight modes, such as Takeoff and Landing. While the parameter ranges provided by UAV configuration manuals are useful, they typically represent general or broader scenarios. Testers often focus on evaluating UAV application-specific or context-dependent conditions. For example, a tester might need to assess parameters like the UAV’s flight range, position, and navigation in a crop monitoring context. Similarly, in a package delivery scenario, a tester could be interested in evaluating parameters such as maximum weight, speed, and battery life along the intended flight path. For testing such conditions, testers must define OCL constraints considering each testing scenario. In our approach, these constraints are utilized to analyze deviations from the expected behavior of the SUT without imposing any restrictions on UAV flight operations. Therefore, OCL constraints must be carefully modeled to capture the testable range of values for the various properties during different flight states. It is important to note that each UAV has its unique set of constraints. The constraints defined for one specific UAV may not necessarily apply to other models of UAVs.
LABEL:lst:constraints presents a set of example constraints designed to demonstrate the process of constraint specification, without being specific to any UAV type. Constraints C1 and C2 define general value ranges for altitude_AGL and distance properties of a UAV. These constraints include ranges that could be obtained from a UAV’s configuration manual. Constraint C1 specifies that a UAV can reach up to a maximum altitude of 300m above ground level (AGL). Exceeding this limit indicates a deviation from the expected behavior, suggesting a potential malfunction in the SUT. Similarly, constraint C2 defines a range for distance measurement by the UAV’s rangefinder, which is particularly beneficial for avoiding potential obstacles. Constraints C3 and C4 specify the valid ranges for the thrust and altitude_AGL during the Takeoff state. In the same way, constraints C5 and C6 define the valid range of values for the airspeed and speed along the z-axis during the Landing state. The constraints for the Takeoff and Landing states can be derived from the UAV’s configuration manual, as it defines the normal ranges for critical flight modes. The takeoff state is crucial for initiating a stable flight for the UAV, while the landing state is essential for the UAV’s safe return to the ground. For instance, a UAV surpassing the takeoff altitude limit could lose control, and violating landing speeds may risk a crash. Constraints C7 and C8 define ranges for airspeed (i.e., between 10 and 100 knots) and groundspeed (i.e., between 0 and 10 knots) during the Ascend state. With these constraints, a tester would be interested in examining the safety scenarios related to the UAV’s airspeed and groundspeed during the ascend/climb phase. For example, maintaining a low airspeed might risk stalling by compromising the UAV’s lift and a high groundspeed could destabilize the UAV. Constraints C9 and C10 specify valid ranges for airspeed (specifically between 5 and 100 knots) and altitude_AGL (specifically between 10 and 100 meter) during the Descend state. These constraints are necessary for UAV’s descending phase and a tester might be interested in testing scenarios related to this phase. For instance, a high airspeed or a low altitude could potentially result in a UAV crash.
@@ 3.4 System-level Testing of UAS
A typical practice of the avionics tester is to simulate the flight behavior of the UAV using the SIL/HIL simulator. During the simulation, the tester performs a certain action, observes the next state of the UAV in an environment, and analyzes the flight data to see if each avionics component is behaving according to the specification. This process of testing requires control of the UAV, continuously deciding various actions, and observing state change in the UAV and environment. To automate this process, human-level decision-making and control of the UAV are required at runtime. Such a problem can be devised in the form of a Markov Decision Process (MDP) [23]. In this case, an AI technique, deep reinforcement learning (DRL) is an appropriate choice compared to meta-heuristics search algorithms. Moreover, DRL provides a mechanism for on-the-fly testing of UAS in a dynamic environment which cannot be achieved by traditional model-based testing approaches (e.g., [24]). Using DRL, our approach introduces an AITester in place of a human tester to perform the system-level testing of the UAV. We formulate this problem as an MDP which consists of a set of states, a set of possible actions that can be performed in a state, a transition model, and a reward function [23]. For this purpose, the proposed profiles, UAS structural profile, and UAV behavioral profile play an important role. More discussion on the adaptation of DRL to automate the system-level testing of the UAV is given in the following subsections.
@@ 3.4.1 Definitions
Definition 1 (State Space). A state space is defined as an ordered set of states $S=\{S_{0},S_{1},S_{2},\dots,S_{g}\}$ in which each state is represented by a tuple $S_{x}={\left\langle s_{0},s_{1},\dots,s_{9}\right\rangle}$. Where $S_{g}$ represents the goal (final) state and $s_{0}$ represents one of the flight state stereotypes modeled in the UAV flight state machine. The remaining state elements are denoted as $s_{1}:=altitude$, $s_{2}:=airspeed$, $s_{3}:=groundspeed$, $s_{4}:=roll$, $s_{5}:=pitch$, $s_{6}:=yaw$, $s_{7}:=heading$, $s_{8}:=battery$, and $s_{9}:=distance$. Given the partially observable environment of UAVs [25], we consider the state space as a subset of the observation space. Hence, our definition of the state space incorporates a selection of properties from the UAS profile (Fig. 3), which constitutes the observation space (as per Definition 5). This state space design is based on key properties identified through various experimental trials. We keep the state space size small to avoid the “curse of dimensionality”—a phenomenon where the amount of data required to learn grows exponentially with the state space size, potentially hindering the agent’s learning capability [26]. The main objective is to streamline learning and decision-making processes within high-dimensional environments like UAVs [27].
Fig. 9 shows an example of the different states in the state space considering the UAV flight state machine shown in Fig. 7. For the given example, the initial state is represented as $S_{0}={\left\langle\mbox{\em Disarmed},0.0,0.0,0,0,0.0,0.0,0.0,1,100,0.0\right\rangle}$, where Disarmed corresponds to $s_{0}$ in the tuple $S_{x}$. In this state, all the values are zero except the battery level which is 100. The next state is $S_{1}={\left\langle\mbox{\em Armed},0.0,0.05,0.2,0.0081,0.0084,0.0235,357,99,0.0\right\rangle}$ where the airspeed, groundspeed, roll, pitch, yaw, heading, and battery values are little changed. After the UAV is armed, the next flight state is Takeoff. Therefore, the next state is $S_{2}={\left\langle\mbox{\em Takeoff},20.0,0.22,2.5,-0.004,-0.004,-0.0501,357,98,0.0\right\rangle}$ in which the altitude is increased to 20.0m, the airspeed is increased to 0.22, the groundspeed is increased to 2.5, and there is a small change in roll, pitch, yaw, and battery values. Similarly, the UAV state continues to change i.e., from $S_{2}$ to $S_{3}$ to $S_{4}$ to $S_{5}$ to $S_{6}$ to $S_{7}$ state which is the goal (final) state. The goal state $S_{g}$ is represented as $S_{7}={\left\langle\mbox{\em Disarmed},0.0,0.0,0.0,0.0031,0.00039,0.0504,357,90,20.0\right\rangle}$.
@@
Definition 5 (Observation Space). An observation space is defined as a vector of UAV properties, i.e., $O=[o_{0},o_{1},o_{2},\dots,o_{n}]$ from the UAS SUT domain model, created using the UAS structural profile. Where $o_{x}$ corresponds to the UAV properties such as thrust, airspeed, vx, vy, and vz.
@@ 3.4.2 Reward Function
The objective of AITester is to take control of the UAV flight during simulation, perform various actions on the UAV, and maximize the number of violations of the expected behavior. The reward function is designed based on two parameters, (i) correct/incorrect actions, and (ii) violations of the expected behavior specified as OCL constraints. These two parameters are important because a series of correct/incorrect actions may lead a UAV to crash. For example, if AITester continues to execute a correct action decreaseAlt() multiple times in Descent state, the UAV can hit the ground leading it to crash. This can happen frequently during training because the inherent behavior of the DRL algorithm involves exploring various possibilities of actions and observing their outcomes. Therefore, a correct action leading the UAV to crash is not considered a fault in the UAV. To handle such a situation, our approach utilizes the deviations from a UAV’s expected behavior based on violations of OCL constraints. The trace of OCL violations (i.e., deviations from the expected behavior) can support diagnosing faults in the SUT. Since OCL constraints are specified corresponding to each flight state of the UAV, it is possible to have multiple violations of OCL constraints at different states during the whole flight.
@@
$r=\left\{\begin{array}[]{ l l }1+m&\quad correct\ \ action\\ -1&\quad incorrect\ \ action\end{array}\right.$ (2)
@@
The cumulative reward for a single episode is calculated using Equation 3. According to this Equation, the cumulative reward is the added sum of the reward for individual action performed at time t (calculated using Equation 2) multiplied with discount factor ${\gamma}^{t}$ at time step t. The discount factor $\gamma$ is between 0 and 1, i.e., $\gamma\in[0,1]$. This factor is essential in determining the extent to which the AITester prefers future rewards over immediate ones. If the discount factor is closer to 1, the AITester gives more preference to future rewards. This means the AITester will explore more of the state space, even if immediate rewards are not apparent. This can lead to broader state space coverage as the AITester is incentivized to explore and comprehend as much of the environment as possible. On the other hand, if the discount factor is closer to 0, the AITester is more inclined towards immediate rewards. This can limit the exploration of the state space. As a result, the AITester may not thoroughly explore the environment, possibly missing out on regions of the state space that could provide higher long-term rewards. Therefore, to enable broader state space coverage and thorough environment exploration, the discount factor should ideally be adjusted closer to 1, aligning with similar practices in the domain [28]. This value will progressively decrease after each time step t, thereby systematically shifting AITester’s focus toward immediate rewards. Consequently, the AITester initially prefers future rewards and gradually transitions its focus towards immediate rewards as time progresses.
3.4.3 Determining LSTM Architecture for AITester
The UAV flight data at a particular time is correlated to the flight data at the previous instance of time. Due to this, UAV flight data holds a temporal relationship based on environmental context which is highly dynamic. Retaining information among these data points in subsequent time steps is important to comprehend the flight context. Moreover, our approach utilizes Deep Q-Network (DQN) [28] which works on the experience replay method by taking random samples from the replay memory of past experiences. This is an inherent feature of DQN to make learning efficient, however, it breaks correlation among different data points which makes it difficult to acquire the flight context. In such a scenario, an LSTM network is a suitable choice due to its ability to handle sequential data and long-term dependencies.
Figure 10: A single layer representation of LSTM architecture for AITester
@@ 4 Implementation
We implemented the proposed approach in a toolset to support the automated system-level testing of the UAV. All main modules of the toolset are packed into two separate packages. The toolset is available at the GitHub repository1. In the following subsections, we discuss every module of the toolset w.r.t. each package.
@@ 4.1.1 Model Loader
First, this module takes the UAS SUT domain model in the form of UML and loads this model in a data structure along with the meta-class elements (i.e., profile stereotypes). Second, it takes the flight behavior model in the form of a UML state machine and traverses the model to extract flight states, events, and transitions. In case the state machine has composite and orthogonal states, these states are flattened using the algorithm presented by Binder [24]. For the modeling manipulation, it uses an external API named Eclipse Modeling Framework (EMF [^3] https://www.eclipse.org/modeling/emf/) that allows it to load and read the model information conveniently. The third function of this module is to load the OCL constraints and state invariants.
@@ 4.1.4 Constraints Evaluator
The core functionality of this module is to evaluate the OCL constraints and state invariants on the instance model containing the UAV flight information. For this purpose, the Eclipse OCL [^4] https://projects.eclipse.org/projects/modeling.mdt.ocl evaluator is used. Before starting the evaluation process, the OCL constraints and state invariants are extracted (using Model Loader) corresponding to each state of the UAV flight. During execution, the flight data received from the environment is sent to Instance Populator. The instance model is used to prepare an OCL evaluator environment for the evaluation of OCL constraints [32, 33]. Before starting the evaluation process, OCL constraints are loaded from a single file corresponding to each state of the UAV flight. This file contains general constraints and state-specific constraints, resembling the format depicted in LABEL:lst:constraints. All OCL constraints and state invariants are evaluated on the populated instance model. If an OCL constraint or state invariant fails on the instance model, the Eclipse OCL evaluator returns false. If the OCL constraint or state invariant is passed on the instance model, the Eclipse OCL evaluator returns true. At the end of the constraints evaluation, the evaluator returns the total number of failed OCL constraints (violations in the expected behavior of the SUT) for the given flight data.
@@ 4.2 AITester
This constitutes the primary component of the toolset. The key functions of this component are UAV environment preparation, interaction with Constraints Evaluator for the reward calculation, and training and evaluation modes for system-level testing of UAVs. Since AITester is aimed to be used at different levels of testing, the tool contains a training module as well as an evaluation module. All modules of the AITester are developed using Python programming language. The deep learning framework used for this purpose is PyTorch [34].
@@ 5 Empirical Evaluation
This section presents the empirical evaluation of AITester including two experiments. The first experiment was performed using the ArduCopter autopilot system and the second experiment was conducted using an industrial case study of ground control station (GCS) cockpit display systems (CDS), i.e., GCS-CDS. In the following subsections, we discuss research questions, case studies, both experiments’ setups, the design and execution of the experiments, the results of both experiments, and the possible threats to the validity of experimental results.
@@ 5.1 Research Questions
AITester utilizes a UAV flight state machine for the system-level testing of the UAS system under test (SUT) with the goal of violating the expected behavior of the SUT. There are two key aspects for analyzing the performance of the AITester. The first is the effectiveness of AITester in generating test scenarios that can lead to deviations from the expected behavior of the UAS SUT. During the testing process, the second aspect is the ability of the AITester to explore diverse paths from the UAV flight state machine. We formulate the following research questions for the experimental evaluation based on these two aspects.
@@
Both RQs are important from the testing perspective of a UAS SUT. The goal of the AITester is to perform system-level testing of a UAS SUT using a state machine representing the flight behavior of a UAV. RQ1 assesses AITester in terms of violations of the expected behavior of a UAS SUT that can lead to potential faults. This is a primary concern of testing a UAS. Since a flight state machine is utilized in the testing process, it is quite possible that AITester learns to explore a specific set of flight states and actions leading to maximum violations using the same set of OCL constraints. This results in testing the same behavior multiple times which is not a desired feature. Therefore, the analysis of diversity in exploring flight paths is an important consideration.